chad has a total of 5 people in his family.
•Sally is Chad's twin sister.
•Chad's sister Hailey is 4 years older than Chad.
•Chad's mother was 24 when chad was born.
•Chad's father is 3 years older than his mother.
1.Write an expression for each family member's age as a function of Chad's age(x).
2.The average age of the 5 family members is 22. Determine the age of chad's sister Hailey.
Please help me as soon as possible... thank you very much!!!

Answer by 303795(562)

Chad = x
Sally = x
Hailey = x + 4
Mum = 24 + x
Dad = 27 + x
Total family age = 5x + 55
Average = total family age / Number of family members
22 = (5x + 55)/5
22 * 5 = 5x + 55
110 = 5x + 55
110 - 55 = 5x
55 = 5x
x = 11
Chad is 11, Sally is 11, Hailey is 15, Mum is 35, Dad is 38